Ballery1:

(Implicit Diff) can someone please check my work? the answer is 4 but i’m getting 3 for some reason.

Ballery1:

SemiDefinite:

Your calculus looks off. From Product and Chain rules: \(\dfrac{d}{dt} \left(xy = y^4 - x \right) \) gives you \( \dfrac{dx}{dt}y + x \dfrac{dy}{dt} = 4y^3 \dfrac{dy}{dt} - \dfrac{dx}{dt}\) Then plug in the values